OpenStudy (mlgpro):

A triangle has a leg that measures 3 and a hypotenuse of 9. Which leg measure would form a right triangle? twelve ninety square root of seventy-two square root of ninety

OpenStudy (anonymous):

a^2 + b^ 2 = c ^ 2 you already know a which is 3 and you know c which is 9. solve for b

OpenStudy (mlgpro):

how

OpenStudy (mlgpro):

how

OpenStudy (mlgpro):

@tejasvir how

OpenStudy (anonymous):

thats the formula for a right triangle

OpenStudy (mlgpro):

6^2

OpenStudy (anonymous):

hm well 81 - 9 = b^2 solve for b